# Bimetric MOND as a framework for variable-$G$ theories – local systems and cosmology

### v1

@@ Abstract
Bimetric MOND (BIMOND) is used as a platform for variable-$G$ theories that have MOND-specific idiosyncrasies. E.g., MOND premises dictate return to standard dynamics in the high-acceleration limit, predicting the standard value of $G$ for high-acceleration systems. This automatically ensures compliance of such theories with all the constraints on inconstancy of $G$ that emerge from the study of high-acceleration systems: geophysics, solar system, pulsars, supernovae, stellar evolution, emission of gravitational waves, etc. In MOND, constraints deduced from such phenomena have no bearing on possible $G$ variability in cosmology. My guiding motivation is to see if such theories may account for some roles of dark matter in cosmology; e.g., in accounting for the expansion history of the Universe in the matter-dominated era, by having a $G_{e}\approx 2\pi G$ govern the later stages of the expansion, instead of invoking matter density $\approx 2\pi\times$ baryon density. BIMOND is a relativistic MOND theory whose gravitational degrees of freedom are two metrics: $g_{\scriptscriptstyle\mu\nu}$, to which matter couples minimally, and $\hat{g}_{\scriptscriptstyle\mu\nu}$, defining the geometry of a twin sector, presumably with its own “twin-matter”. The BIMOND gravitational action for the metric-symmetric subclass is $I^{\scriptscriptstyle B}\propto G^{-1}\int d^{4}x~[|g|^{\scriptscriptstyle 1/2}R+|\hat{g}|^{\scriptscriptstyle 1/2}\hat{R}+v(g,\hat{g})\ell_{\scriptscriptstyle M}^{\scriptscriptstyle-2}\tilde{\mathcal{M}}$], where the metric-interaction term, $\tilde{\mathcal{M}}$, is a function of dimensionless scalars constructed from $\ell_{\scriptscriptstyle M}C^{\scriptscriptstyle\alpha}_{\scriptscriptstyle\beta\gamma}$, where the “relative-acceleration” tensor, $C^{\scriptscriptstyle\alpha}_{\scriptscriptstyle\beta\gamma}\equiv\Gamma^{\scriptscriptstyle\alpha}_{\scriptscriptstyle\beta\gamma}-\hat{\Gamma}^{\scriptscriptstyle\alpha}_{\scriptscriptstyle\beta\gamma}$, is the difference of the Levi-Civita connections, and $\ell_{\scriptscriptstyle M}=c^{2}/a_{0}$ is the MOND length. Locally, where the random fluctuations in the two metrics are different, $\tilde{\mathcal{M}}$ encapsulates MOND effects in galactic systems, and the average of $\tilde{\mathcal{M}}$ on cosmological scales acts as a (dynamical) dark energy. Without adding degrees of freedom, or new dimensionful constants, BIMOND can be extended to a class of theories that entail what is best described as phenomenon-dependence of Newton’s constant, $G$: The Lagrangian density in $I^{\scriptscriptstyle B}$ is multiplied by $1+\mathcal{M}_{\scriptscriptstyle G}$, where $\mathcal{M}_{\scriptscriptstyle G}$ too is a function of the above scalars. $G_{e}\equiv G/(1+\mathcal{M}_{\scriptscriptstyle G})$ plays the role of the effective gravitational constant. I consider the standard and the Einstein-Palatini formulations for the above action. My aim here is to present the general framework, while specific predictions would depend crucially on details of the structure of the theory – e.g., on the choice of variables, and the dependence of $\tilde{\mathcal{M}}$ and $\mathcal{M}_{\scriptscriptstyle G}$ on them – and, in cosmology, on unknown inputs: cosmological initial conditions, material content of the twin sector, etc. I cannot yet present a consistent model that complies with all the observations in cosmology, including the expansion history, with all its details. Instead, I describe some examples of theories in the class that predict different values of $G_{e}$ in different circumstances, including one where $G$ takes its standard value for all subcosmological systems – even if they are deep in the MOND regime. I also discuss scenarios in which $G_{e}\approx G$ in the early Universe, as required by constraints from big-bang nucleosynthesis, but with $G_{e}>G$ setting in at later times, where it can affect the expansion history during the matter-dominated era.
@@ I Introduction
The exact structure-formation scenario in MOND is still moot. But several works have shown [12, 13, 14, 15, 16] that it occurs, generically, in a different way than in the $\Lambda{\rm CDM~}$paradigm, with galaxies and larger-scale structure forming rather earlier, as, in fact, appears to be vindicated by recent observations [16]. There are also natural explanations for the appearance of a cosmological constant, or “dark energy”, within various MOND theories ([11], and Sec. VI below).
Putting aside, then, the important question of what replaces (in MOND) the role of dark matter in CMB fluctuations, and structure formation, we ask, as a limited question, whether MOND can account, not only for the role of dark energy, but also for that of dark matter, in dictating the observed expansion history of the Universe. A possible pointer towards effecting this end is the fact that in this context, the densities of (NR) baryons, $\rho_{b}$, and of DM, $\rho_{\scriptscriptstyle DM}$, enter in the combination $G(\rho_{b}+\rho_{\scriptscriptstyle DM})$, and, to account for the expansion history with dark matter it is required that $\rho_{\scriptscriptstyle M}=\rho_{b}+\rho_{\scriptscriptstyle DM}\approx 2\pi\rho_{b}$ (e.g., Ref. [17]). Thus, the same history results, without dark matter, if the effective $G$ value that governs the expansion history during the relevant epoch is $G_{\scriptscriptstyle e}\approx 2\pi G$.
@@
Predictions of such theories cannot, generally, be described as those of standard dynamics only with a different, spacetime-dependent, [^1] As always, variability of $G$ is understood as the variability of its value as measured in atomic units. effective $G_{e}=G/f$. But, such an ansatz is sometimes justified approximately when testing and constraining the possible inconstancy of $G$: One does, usually, consider some phenomenon – e.g., big-bang nucleosynthesis, stellar evolution, pulsar timing, etc. – and analyzes it under the above ansatz. A comprehensive review of such tests can be found in Ref. [18] (Sec. 6 there is dedicated to tests of, and limits on, possible inconstancy of $G$).
@@
Since MOND introduces a new dimensionful “dividing constant”, [^2] By a “dividing constant”, I mean a physical constant, $\mathcal{C}$, that, among other roles, defines a dividing line: Systems whose attributes, $\mathcal{Y}$, with the dimensions of $\mathcal{C}$, behave differently – e.g., obey different scaling relations – depending on whether $\mathcal{Y}/\mathcal{C}$ is small or large. Usually, $\mathcal{C}$ itself appears ubiquitously in relations that hold in one of the limits, and disappear in those valid in the other limit. This clearly include $c$ in relativity, $\hbar$ in quantum physics, and $a_{0}$ in MOND, but not $G$, or Boltzmann’s constant, which are only “conversion” constants, and do not play such a role. $a_{0}$, together with $c$ they define a fourfold classification of phenomena and systems, according to whether they are relativistic or not, and according to whether they are high-acceleration or not, in the MOND sense. All subcosmological phenomena for which tight constraints on $G$ inconstancy exist, are either NR, or characterized by very-high-accelerations, because the “coincidence” (1) tells us that relativistic, subcosmological systems are perforce of high accelerations. This puts all of them in separate classes from cosmology. This makes it natural in MOND to have $G$-inconstancy effects in cosmology, without them showing up in subcosmological systems. [^3] When dealing with a VGT that modifies GR (and retains the standard Newtonian limit), but does not introduce new dimensionful constants – as is the case in the iconic Jordan-Brans-Dicke VGT [19] – we have a twofold classification of systems according to whether they are relativistic or not. The strength of $G$-inconstancy effects then depends on how relativistic the system is, which puts cosmology in the same class as black holes, pulsar phenomena, compact-stars mergers with the emission of gravitational waves, etc. Then, tight constraints obtained from relativistic, subcosmological systems and phenomena, bear, directly, on $G$ inconstancy in cosmology. Such an isolation of cosmology as the only phenomenon in the strong-gravity-low-acceleration quadrant would underlie any VGT theory based on MOND. Here I shall demonstrate it specifically with BIMOND-based VGTs. BIMOND-based VGTs may have additional idiosyncrasies that bear on the issue, and that are not necessarily shared by other MOND-based VGTs.
@@
I do not demonstrate here that there are theories in this class that conforms to all constraints from cosmology. At present, the suggested class of theories remains an interesting framework for VGTs that aligns with MOND, from which they borrow heavily.
@@
In BIMOND, specifically, there is an additional idiosyncrasy that affords an interesting leverage: It is that in situation of equality of the two metrics, all the scalar variable vanish, and, more generally, the departure from equality is a parameter that can control the departure of $G_{e}$ from $G$. I use this fact to suggest how in a cosmology that starts with an exactly metric-symmetric initial conditions, initially, $G_{e}=G$, which would comply with the constraints from big-bang nucleosynthesis, with a departure taking place only at later epochs, when asymmetric random fluctuations develop and increase.
@@
I end this introduction by a general comment on relativistic MOND theories. Inasmuch as they tend quickly to GR for high-acceleration systems, such theories are not required to account for relativistic objects whose dynamics depart from GR, except for cosmology, which is a unique, and uncontrollable system. I think then that we may be more forgiving to such theories of some issues that might only appear in applications to local systems (such as Ostrogradky’s instabilities, or the appearance of ghosts (see Ref. [30] for a discussion of such issues with the first version of BIMOND). For such systems, the theory practically coincides with GR, which is free of these issues. And if such theories turn out not to be “fundamental” – which known working theory is? – they can at least serve as useful heuristic tools. In the present discussion, they can limelight the idiosyncracies of MOND based VGTs.
@@ II BIMOND recap
for which the two sectors decouple for interesting cases where $\tilde{\mathcal{M}}$ is constant. The matter actions, not shown here, are the standard ones with matter in each sector coupling minimally to its own metric. [^5] More general versions may be, and have been, considered; e.g., with lack of symmetry between the two sectors, with different volume elements, with non-quadratic scalars, etc.
@@
The “correspondence principle” whereby the theory goes to GR in the limit of high acceleration – formally by taking $\mathcal{Z}\gg 1$ is ensured by having $\tilde{\mathcal{M}}$ becoming a (dimensionless) constant, $\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{M}}_{\infty}$, in this limit. Then the two sectors decouple [with the choice (5) of the volume element] and each is governed by GR gravity, with a cosmological constant $\Lambda=-\ell_{\scriptscriptstyle M}^{-2}\tilde{\mathcal{M}}_{\infty}/2$.
@@ II.1 The BIMOND scalar variables
Here I describe the scalar variables, their behaviour in various relevant circumstances, and constraints on their choice that we already have from observed phenomenology. These are important to appreciate when constructing VGMOND  theories.
@@ II.1.1 Lensing constraints on the scalars
Because there are two forms of mixed terms, $\phi^{\scriptscriptstyle*}_{\scriptscriptstyle,i}h^{\scriptscriptstyle*}_{\scriptscriptstyle,i}$ and $\phi^{\scriptscriptstyle*}_{\scriptscriptstyle,i}h^{\scriptscriptstyle*}_{\scriptscriptstyle ik,k}$, eliminating them leaves us with a three-parameter class of scalars – call them “good scalar of the first type” – whose general form is [24]
@@ III.1 Field equation
While this is a “higher-derivative theory” – such as would generically (but not always) lead to the appearance of Ostrogradsky instabilities [32] – note that it is a degenerate one, in the sense that is relevant for such theories, since the second derivatives appear linearly in the Lagrangian density. Such degenerate theories are not necessarily beset by Ostrogradsky instabilities. But this needs to be checked specifically the present case for the variety of possible theories in the class. Degenerate, higher-derivative theories, with a function of a scalar field and its first derivative appearing as a prefactor of the Ricci scalar (instead of $1+\mathcal{M}_{\scriptscriptstyle G}$ here) have been discussed, e.g., in Refs. [33, 34].
@@
If indeed it would dominate, such a term would greatly depart from the NR limit of BIMOND itself. However, for high-acceleration systems, the MOND tenets require that $\mathcal{M}_{\scriptscriptstyle G}\lll 1$, and it is easy to make this term strongly subdominant. But, in order to preserve the performance of BIMOND also in low-acceleration systems, such as galaxies, we should also require that $\mathcal{M}_{\scriptscriptstyle G}$ vanishes rapidly in the NR limit for all accelerations. This can be done, as I show in Sec. V.0.4, below.
I shall not further expand expressions (26) and (27), the resulting, cumbersome expressions can be straightforwardly written using formulae in Ref. [24] (Sec. IIIA there), replacing there $\tilde{\mathcal{M}}$ by $\mathcal{M}_{\scriptscriptstyle G}$ – the variable on which these functions depend coming from the same stock.

### v2

# Bimetric MOND as a framework for variable-$G$ theories: Local systems and cosmology

@@ Abstract
Bimetric MOND (BIMOND) is used as a platform for variable-$G$ theories that have MOND-specific idiosyncrasies. E.g., MOND premises dictate return to standard dynamics in the high-acceleration limit, predicting the standard value of $G$ for high-acceleration systems. This automatically ensures compliance of such theories with all the constraints on inconstancy of $G$ that emerge from the study of high-acceleration systems: geophysics, solar system, pulsars, supernovae, stellar evolution, emission of gravitational waves, etc. In MOND, constraints deduced from such phenomena have no bearing on possible $G$ variability in cosmology. My guiding motivation is to see if such theories may account for some roles of dark matter in cosmology; e.g., in accounting for the expansion history of the Universe in the matter-dominated era, by having a $G_{e}\approx 2\pi G$ govern the later stages of the expansion, instead of invoking matter density $\approx 2\pi\times$ baryon density. BIMOND is a relativistic MOND theory whose gravitational degrees of freedom are two metrics: $g_{\scriptscriptstyle\mu\nu}$, to which matter couples minimally, and $\hat{g}_{\scriptscriptstyle\mu\nu}$, defining the geometry of a twin sector, presumably with its own “twin-matter.” The BIMOND gravitational action for the metric-symmetric subclass is $I^{\scriptscriptstyle B}\propto G^{-1}\int d^{4}x~[|g|^{\scriptscriptstyle 1/2}R+|\hat{g}|^{\scriptscriptstyle 1/2}\hat{R}+v(g,\hat{g})\ell_{\scriptscriptstyle M}^{\scriptscriptstyle-2}\tilde{\mathcal{M}}$], where the metric-interaction term, $\tilde{\mathcal{M}}$, is a function of dimensionless scalars constructed from $\ell_{\scriptscriptstyle M}C^{\scriptscriptstyle\alpha}_{\scriptscriptstyle\beta\gamma}$, where the “relative-acceleration” tensor, $C^{\scriptscriptstyle\alpha}_{\scriptscriptstyle\beta\gamma}\equiv\Gamma^{\scriptscriptstyle\alpha}_{\scriptscriptstyle\beta\gamma}-\hat{\Gamma}^{\scriptscriptstyle\alpha}_{\scriptscriptstyle\beta\gamma}$, is the difference of the Levi-Civita connections, and $\ell_{\scriptscriptstyle M}=c^{2}/a_{0}$ is the MOND length. Locally, where the random fluctuations in the two metrics are different, $\tilde{\mathcal{M}}$ encapsulates MOND effects in galactic systems, and the average of $\tilde{\mathcal{M}}$ on cosmological scales acts as a (dynamical) dark energy. Without adding degrees of freedom, or new dimensionful constants, BIMOND can be extended to a class of theories that entail what is best described as phenomenon-dependence of Newton’s constant, $G$: The Lagrangian density in $I^{\scriptscriptstyle B}$ is multiplied by $1+\mathcal{M}_{\scriptscriptstyle G}$, where $\mathcal{M}_{\scriptscriptstyle G}$ too is a function of the above scalars. $G_{e}\equiv G/(1+\mathcal{M}_{\scriptscriptstyle G})$ plays the role of the effective gravitational constant. I consider the standard and the Einstein-Palatini formulations for the above action. My aim here is to present the general framework, while specific predictions would depend crucially on details of the structure of the theory – e.g., on the choice of variables, and the dependence of $\tilde{\mathcal{M}}$ and $\mathcal{M}_{\scriptscriptstyle G}$ on them – and, in cosmology, on unknown inputs: cosmological initial conditions, material content of the twin sector, etc. I cannot yet present a consistent model that complies with all the observations in cosmology, including the expansion history, with all its details. Instead, I describe some examples of theories in the class that predict different values of $G_{e}$ in different circumstances, including one where $G$ takes its standard value for all subcosmological systems – even if they are deep in the MOND regime. I also discuss scenarios in which $G_{e}\approx G$ in the early Universe, as required by constraints from big bang nucleosynthesis, but with $G_{e}>G$ setting in at later times, where it can affect the expansion history during the matter-dominated era.
@@ I Introduction
The exact structure-formation scenario in MOND is still moot. But several works have shown [12, 13, 14, 15, 16] that it occurs, generically, in a different way than in the $\Lambda{\rm CDM~}$paradigm, with galaxies and larger-scale structure forming rather earlier, as, in fact, appears to be vindicated by recent observations [16].
There are also natural explanations for the appearance of a cosmological constant, or “dark energy”, within various MOND theories ([11], and Sec. VI below).
Putting aside, then, the important question of what replaces (in MOND) the role of dark matter in CMB fluctuations, and structure formation, we ask, as a limited question, whether MOND can account, not only for the role of dark energy, but also for that of dark matter, in dictating the observed expansion history of the Universe. A possible pointer toward effecting this end is the fact that in this context, the densities of (NR) baryons, $\rho_{b}$, and of DM, $\rho_{\scriptscriptstyle DM}$, enter in the combination $G(\rho_{b}+\rho_{\scriptscriptstyle DM})$, and, to account for the expansion history with dark matter it is required that $\rho_{\scriptscriptstyle M}=\rho_{b}+\rho_{\scriptscriptstyle DM}\approx 2\pi\rho_{b}$ (e.g., Ref. [17]). Thus, the same history results, without dark matter, if the effective $G$ value that governs the expansion history during the relevant epoch is $G_{\scriptscriptstyle e}\approx 2\pi G$.
@@
Predictions of such theories cannot, generally, be described as those of standard dynamics only with a different, spacetime-dependent, [^1] As always, variability of $G$ is understood as the variability of its value as measured in atomic units. effective $G_{e}=G/f$. But, such an ansatz is sometimes justified approximately when testing and constraining the possible inconstancy of $G$: One does, usually, consider some phenomenon – e.g., big bang nucleosynthesis, stellar evolution, pulsar timing, etc. – and analyzes it under the above ansatz. A comprehensive review of such tests can be found in Ref. [18] (Sec. 6 there is dedicated to tests of, and limits on, possible inconstancy of $G$).
@@
Since MOND introduces a new dimensionful “dividing constant”, [^2] By a “dividing constant,” I mean a physical constant, $\mathcal{C}$, that, among other roles, defines a dividing line: Systems whose attributes, $\mathcal{Y}$, with the dimensions of $\mathcal{C}$, behave differently – e.g., obey different scaling relations – depending on whether $\mathcal{Y}/\mathcal{C}$ is small or large. Usually, $\mathcal{C}$ itself appears ubiquitously in relations that hold in one of the limits, and disappears in those valid in the other limit. This clearly include $c$ in relativity, $\hbar$ in quantum physics, and $a_{0}$ in MOND, but not $G$, or Boltzmann’s constant, which are only “conversion” constants, and do not play such a role. $a_{0}$, together with $c$ they define a fourfold classification of phenomena and systems, according to whether they are relativistic or not, and according to whether they are high-acceleration or not, in the MOND context. All subcosmological phenomena for which tight constraints on $G$ inconstancy exist, are either NR, or characterized by very-high-accelerations, because the “coincidence” (1) tells us that relativistic, subcosmological systems are perforce of high accelerations. This puts all of them in separate classes from cosmology. This makes it natural in MOND to have $G$-inconstancy effects in cosmology, without them showing up in subcosmological systems. [^3] When dealing with a VGT that modifies GR (and retains the standard Newtonian limit), but does not introduce new dimensionful constants – as is the case in the iconic Jordan-Brans-Dicke VGT [19] – we have a twofold classification of systems according to whether they are relativistic or not. The strength of $G$-inconstancy effects then depends on how relativistic the system is, which puts cosmology in the same class as black holes, pulsar phenomena, compact-stars mergers with the emission of gravitational waves, etc. Then, tight constraints obtained from relativistic, subcosmological systems and phenomena, bear, directly, on $G$ inconstancy in cosmology. Such an isolation of cosmology as the only phenomenon in the strong-gravity-low-acceleration quadrant would underlie any VGT theory based on MOND. Here I shall demonstrate it specifically with BIMOND-based VGTs. Such theories may have additional idiosyncrasies that bear on the issue, and that are not necessarily shared by other MOND-based VGTs.
@@
I do not demonstrate here that there are theories in this class that conform to all constraints from cosmology. At present, the suggested class of theories remains an interesting framework for VGTs that aligns with MOND, from which they borrow heavily.
@@
In BIMOND, specifically, there is an additional idiosyncrasy that affords an interesting leverage: It is that in situation of equality of the two metrics, all the scalar variable vanish, and, more generally, the departure from equality is a parameter that can control the departure of $G_{e}$ from $G$. I use this fact to suggest how in a cosmology that starts with an exactly metric-symmetric initial conditions, initially, $G_{e}=G$, which would comply with the constraints from big bang nucleosynthesis, with a departure taking place only at later epochs, when asymmetric random fluctuations develop and increase.
Note, in this connection, that the accounts of galaxy formation in Refs. [12, 13, 14, 15, 16], and others, are not based on some relativistic theory that takes into account the cosmological expansion, together with the gravitational collapse of structure. In particular, they all assume that the acceleration that determines the collapse dynamics is only that contributed by the local overdensity, and ignore the background acceleration associated with the expansion. Metric-symmetric versions of BIMOND justify this ansatz formally. What enter their MOND aspects that govern the dynamics of local systems are the “relative acceleration” scalars, in which the background expansion cancels out (see Sec. VI).
@@
I end this introduction by a general comment on relativistic MOND theories. Inasmuch as they tend quickly to GR for high-acceleration systems, such theories are not required to account for relativistic objects whose dynamics depart from GR, except for cosmology, which is a unique system. I think then that we may be more forgiving to such theories of some issues that might only appear in applications to local systems (such as Ostrogradky’s instabilities, or the appearance of ghosts (see Ref. [30] for a discussion of such issues with the first version of BIMOND). For such systems, the theory practically coincides with GR, which is free of these issues. And if such theories turn out not to be “fundamental” – which known working theory is? – they can at least serve as useful heuristic tools. In the present discussion, they can limelight the idiosyncracies of MOND based VGTs.
@@ II BIMOND recap
for which the two sectors decouple for interesting cases where $\tilde{\mathcal{M}}$ is constant. The matter actions, not shown here, are the standard ones with matter in each sector coupling minimally to its own metric. [^5] More general versions may be, and have been, considered; e.g., with lack of symmetry between the two sectors, with different volume elements, with nonquadratic scalars, etc.
@@
The “correspondence principle” whereby the theory goes to GR in the limit of high acceleration – formally by taking $\mathcal{Z}\gg 1$ – is ensured by having $\tilde{\mathcal{M}}$ becoming a (dimensionless) constant, $\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{M}}_{\infty}$, in this limit. Then the two sectors decouple [with the choice (5) of the volume element] and each is governed by GR gravity, with a cosmological constant $\Lambda=-\ell_{\scriptscriptstyle M}^{-2}\tilde{\mathcal{M}}_{\infty}/2$.
@@
Another class of bimetric extensions of GR that were proposed to account for MOND phenomenology are described in Ref. [31].
@@ II.1 The BIMOND scalar variables
Here I describe the scalar variables, their behavior in various relevant circumstances, and constraints on their choice that we already have from observed phenomenology. These are important to appreciate when constructing VGMOND  theories.
@@ II.1.1 Lensing constraints on the scalars
Because there are two forms of mixed terms, $\phi^{\scriptscriptstyle*}_{\scriptscriptstyle,i}h^{\scriptscriptstyle*}_{\scriptscriptstyle,i}$ and $\phi^{\scriptscriptstyle*}_{\scriptscriptstyle,i}h^{\scriptscriptstyle*}_{\scriptscriptstyle ik,k}$, eliminating them leaves us with a three-parameter class of scalars – call them “good scalars of the first type” – whose general form is [24]
@@ III.1 Field equation
While this is a higher-derivative theory – such as would generically (but not always) lead to the appearance of Ostrogradsky instabilities [33] – note that it is a degenerate one, in the sense that is relevant for such theories, since the second derivatives appear linearly in the Lagrangian density. Such degenerate theories are not necessarily beset by Ostrogradsky instabilities. But this needs to be checked specifically the present case for the variety of possible theories in the class. Degenerate, higher-derivative theories, with a function of a scalar field and its first derivative appearing as a prefactor of the Ricci scalar (instead of $1+\mathcal{M}_{\scriptscriptstyle G}$ here) have been discussed, e.g., in Refs. [34, 35].
@@
If indeed it would dominate, such a term would greatly depart from the NR limit of BIMOND itself. However, for high-acceleration systems, the MOND tenets require that $\mathcal{M}_{\scriptscriptstyle G}\lll 1$, and it is easy to make this term strongly subdominant. But, in order to preserve the performance of BIMOND also in low-acceleration systems, such as galaxies, we should also require that $\mathcal{M}_{\scriptscriptstyle G}$ vanishes rapidly in the NR limit for all accelerations. This can be done, as I show in Sec. V.4, below.
I shall not further expand expressions (26) and (27), the resulting, cumbersome expressions can be straightforwardly written using formulas in Ref. [24] (Sec. IIIA there), replacing there $\tilde{\mathcal{M}}$ by $\mathcal{M}_{\scriptscriptstyle G}$ – the variable on which these functions depend coming from the same stock.